OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2). (Points : 2) y = 5/2 x - 2 y = -5/2 x + 8 Y = 2/5 x - 6 y = -2/5 x + 4

OpenStudy (anonymous):

m = (8--2)/(4-0) = 10/4 = 5/2. y = (5/2)x + b solve for b by saying = (5/2)(4) + b.